\begin{thebibliography}{0}

\bibitem{oleari_dis06}
C.~Oleari, these proceedings.

\bibitem{aleph_ff}
A.~Heister {\it et al.}, ALEPH Collab.,
{\it Phys. Lett.} {\bf B512}, 30 (2001).

\bibitem{opal_ff}
G.~Abbiendi {\it et al.}, OPAL Collab.,
{\it Eur. Phys. J.} {\bf C29}, 463 (2003).

\bibitem{delphi_ff}
DELPHI Collab.,
%{\it Measurement of charm fragmentation function in $D^*$ photoproduction at HERA}. 
Abstract 583,
International Conference on High Energy Physics,
Amsterdam, The Netherlands (ICHEP 2002), July 2002;\\
http://delphiwww.cern.ch/pubxx/conferences/amsterdam02/.

\bibitem{sld_ff}
K.~Abe {\it et al.}, SLD Collab.,
{\it Phys. Rev.} {\bf D65}, 092006 (2002).
 
\bibitem{pythia}
T.~Sj\"{o}strand,
{\it Comp. Phys. Comm.} {\bf 82}, 74 (1994).
 
\bibitem{kartve}
V.G~Kartvelishvili, A.K.~Likhoded, V.A.~Petrov,
{\it Phys. Lett.} {\bf B78}, 615 (1978).

\bibitem{bowler}
M.G.~Bowler,
{\it Z. Phys.} {\bf C11} 169 (1981).
%X.~Artru, G.~Mennessier, {\it Nucl. Phys.} {\bf B70} 93 (1974).

\bibitem{lund}
B.~Anderson, G.~Gustafson, B.~S\"{o}deberger,
{\it Z. Phys.} {\bf C20} 317 (1983).
 
\bibitem{peterson}
C.~Peterson {\it et al.},
{\it Phys. Rev.} {\bf D27} 105 (1983).
 
\bibitem{collins}
P.D.B.~Collins, T.P.~Spiller,
{\it J. Phys.} {\bf G11} 1289 (1985).
 
\bibitem{herwig}
G.~Marchesini {\it et al.},
{\it Comp. Phys. Comm.} {\bf 67}, 465 (1992).

\bibitem{cacciari_frag}
M.~Cacciari, P.~Nason, C.~Oleari,
{\it JHEP} {\bf 0604}, 006 (2006).

\bibitem{kniehl_bfrag}
J.~Binnewies, B.A.~Kniehl, G.~Kramer,
{\it Phys. Rev.} {\bf D58}, 034016 (1998).

\bibitem{hfag}
Heavy Flavour Averaging Group (HFAG),\\
http://www.slac.stanford.edu/xorg/hfag/osc/.

\bibitem{pdg04}
S.~Eidelman {\it et al.},
{\it Phys. Lett.} {\bf B592}, 1 (2004).

\bibitem{cleo_ff}
M.~Artuso {\it et al.}, CLEO Collab.,
{\it Phys. Rev.} {\bf D70}, 112001 (2004).

\bibitem{belle_ff}
R.~Seuster {\it et al.}, BELLE Collab.,
{\it Phys. Rev.} {\bf D73}, 032002 (2006).

\bibitem{aleph_cc_ff}
R.~Barate {\it et al.}, ALEPH Collab.,
{\it Eur. Phys. J.} {\bf C16}, 597 (2000).

\bibitem{zeus_ffunction}
ZEUS Collab.,
%{\it Measurement of charm fragmentation function in $D^*$ photoproduction at HERA}. 
Abstract 778,
International Conference on High Energy Physics,
Amsterdam, The Netherlands (ICHEP 2002), July 2002;\\
http://www-zeus.desy.de/physics/phch/conf/amsterdam\underline{ }paper.html.

\bibitem{h1_ffunction}
H1 Collab.,
%{\it Measurement of charm fragmentation function in $D^*$ photoproduction at HERA}. 
Abstract 649,
International Europhysics Conference on High Energy Physics,
Lisbon, Portugal (HEP 2005), July 2005;\\
http://www-h1.desy.de/h1/www/publications/conf/list.EPS2005.html.

\bibitem{zeus_php_fb} S.~Chekanov {\it et al.}, ZEUS Collab.,
Eur. Phys. J. {\bf C44} (2005) 351.

\bibitem{h1_dis_fb}
A.~Aktas {\it et al.}, H1 Collab.,
Eur. Phys. J. {\bf C38} (2005) 447.

\bibitem{hep-ex-9912064}
L.~Gladilin, Preprint hep-ex/9912064, 1999.

%\bibitem{lund}
%B.~Anderson {\it et al.}, Phys. Rep. {\bf 97} (1983) 31.

%\bibitem{argus}
%H.~Albrecht {\it et al.}, ARGUS Collab.,
%Z. Phys. {\bf C52} (1991) 353.

%\bibitem{opal}
%R.~Akers {\it et al.}, OPAL Collab.,
%Z. Phys. {\bf C67} (1995) 27.

%\bibitem{oleari}
%P.~Nason and C.~Oleari, Nucl. Phys. {\bf B565} (2000) 245.

%\bibitem{mod_gauss} L.~Gladilin, private communication; available from
%http://www.desy.de/$\sim$gladilin.

%\bibitem{zeus_9697}
%J.~Breitweg {\it et al.}, ZEUS Collab.,
%Phys. Lett. B {\bf481}  (2000) 213.

\bibitem{zeus_dis_fb}
ZEUS Collab.,
%{\it Measurement of charm fragmentation function in $D^*$ photoproduction at HERA}. 
Abstract 351,
International Europhysics Conference on High Energy Physics,
Lisbon, Portugal (HEP 2005), July 2005;\\
http://www-zeus.desy.de/physics/phch/conf/lp05\underline{ }eps05/.

%\bibitem{pdg}
%Particle Data Group, S.~Eidelman {\it et al.},
%Phys. Lett. B {\bf592}  (2004) 1.

%\bibitem{thermo}
%F.~Becattini,
%Z. Phys. {\bf C69} (1996) 485.

%\bibitem{string}
%Yi-Jin~Pei,
%Z. Phys. {\bf C72} (1996) 39.

%\bibitem{bkl}
%A.~V.~Berezhnoy, V.~V.~Kiselev and A.~K.~Likhoded, Phys. Rev. {\bf D62} (2000) 074013;\\
%A.~V.~Berezhnoy and A.~K.~Likhoded, private communication.

\end{thebibliography}

\documentclass{ws-procs9x6}
\newcommand{\fcdz}       {\mbox{$f(c \rightarrow D^0)$}}
\newcommand{\fcdc}       {\mbox{$f(c \rightarrow D^+)$}}
\newcommand{\fcdss}      {\mbox{$f(c \rightarrow D_s^+)$}}
\newcommand{\fclc}       {\mbox{$f(c \rightarrow \Lambda_c^+)$}}
\newcommand{\fcds}       {\mbox{$f(c \rightarrow D^{\ast +})$}}
\newcommand{\dspm}       {\mbox{$D^{\ast \pm}$}}

\begin{document}

\title{Experimental results on heavy quark fragmentation
\footnote{\uppercase{P}articipation in \uppercase{DIS} 2006 was supported by the local organising committee and grant 06-02-26609-z of the
\uppercase{R}ussian \uppercase{F}oundation for \uppercase{B}asic
\uppercase{R}esearch.}}

\author{L.~K.~GLADILIN}

\address{Skobeltsyn Institute of Nuclear Physics, Moscow State University\\
Vorob'evy Gory, Moscow, RU-119992, Russia\\
E-mail: gladilin@sinp.msu.ru }

\maketitle

\abstracts{
Experimental results on $c$- and $b$-quark fragmentation are reviewed.
The discussion is concentrated on measurements of heavy-quark fragmentation
functions
and 
fragmentation fractions.
Measurements of various heavy-quark fragmentation ratios are also discussed.
The experimental results are compared with theoretical expectations and
model predictions.
}

\section{Introduction}

The initial stage of charm/bottom quark fragmentation
can be described by perturbative QCD (pQCD) calculations~\cite{oleari_dis06}.
A non-perturbative (NP) parameterisation is needed to describe
the final heavy-quark transformation to
a particular charmed or bottom hadron.
Such parameterisation can include
effects producing by the excited states decaying to a given hadron.
The NP fragmentation parameterisation can be splited in two
parts: fragmentation function and fragmentation fraction.
Fragmentation functions are used to parameterise the transfer
of the quark's energy to a given meson; they can be different
for different pQCD calculations used to describe the initial
fragmentation.
Fragmentation fractions
are the fractions of $c$/$b$ quarks hadronising as a particular
charmed/bottom hadron; they are expected to be
universal for all pQCD calculations.

Measurements of the heavy quark fragmentation allow testing
pQCD calculations and extracting fragmentation functions
and fractions.
A deeper phenomenological understanding of the heavy quark fragmentation can
be obtained by measuring
various heavy-quark fragmentation ratios.
In particular, we will discuss
the ratio of neutral and charged $D$/$B$ meson production rates, $R_{u/d}$,
the strangeness-suppression factor, $\gamma_s$, and the fraction
of $D$/$B$ mesons produced in a vector state, $P_{\rm v}$.


\section{Bottom quark fragmentation}

The $b$-quark fragmentation function was measured at
LEP~\cite{aleph_ff,opal_ff,delphi_ff}
and SLD~\cite{sld_ff}.
The measured spectra were compared with
predictions of
the leading-logarithmic (LL) JETSET~7.4~\cite{pythia} Monte Carlo (MC)
using different  parametrisations
for the fragmentation function.
The best description of the data with a parametrisation
with one free parameter was obtained using the parametrisation
of Kartvelishvili et al.~\cite{kartve}.
The Bowler~\cite{bowler} and symmetric LUND~\cite{lund}
parametrisations with two free parameters provided a better
data description.
The Peterson~\cite{peterson} and Collins-Spiller~\cite{collins}
parameterisations, and the HERWIG cluster model~\cite{herwig} predictions
were found to be too broad to describe the data.
The $b$-quark fragmentation function measurements were also used for fitting
the NP parametrisation with the next-to-leading-order (NLO)
calculations~\cite{cacciari_frag,kniehl_bfrag}.

The $b$-quark fragmentation fractions
were obtained by combining
of all published LEP and CDF results on production of
the weakly decaying $B$ hadrons with measurements of the time-integrated
mixing probabilities~\cite{hfag,pdg04}.
The isospin invariance, i.e. $R_{u/d}=1$, was assumed in this procedure.
Using the measured fragmentation fractions, the strangeness-suppression factor
for bottom mesons is
$$\gamma_s=2 f(\bar{b}\rightarrow B^0_s)/[f(\bar{b}\rightarrow B^0) + f(\bar{b}\rightarrow B^+)]=0.27\pm0.03.$$
Thus, bottom-strange meson production is suppressed by a factor $\approx3.7$.
The combined LEP value for the fraction
of $B$ mesons, produced in a vector state, is $P_{\rm v}=0.75\pm0.04$~\cite{pdg04}, that is in perfect agreement with
the naive spin counting expectation (0.75).

\section{Charm quark fragmentation}

The $c$-quark fragmentation function has been recently measured with
high precision by the CLEO~\cite{cleo_ff} and
BELLE~\cite{belle_ff} collaborations.
The data comparison with
the JETSET MC predictions revealed the same picture
as for the $b$-quark fragmentation.
The best description of the data was obtained
using the Bowler parametrisation with two free parameters,
and the parametrisation
of Kartvelishvili et al. with one free parameter.

A discrepancy between the NP parametrisations
obtained with
the CLEO/BELLE data and earlier
ALEPH measurement~\cite{aleph_cc_ff}
has been observed
using the NLO initial conditions, next-to-leading
logarithmic (NLL) evolution, NLO coefficient functions and NLL Sudakov
resummation~\cite{oleari_dis06,cacciari_frag}.
The difference, which was attributed
to the evolution between the $\Upsilon(4S)$ and $Z^0$ energies,
results in an additional uncertainty in predictions for $\dspm$
hadroproduction of the order $20\%$.
To reduce the uncertainty 
direct measurements of the charm fragmentation function
at hadronic machines would be useful. Such measurements were
already performed in $ep$ interactions at HERA by
the ZEUS~\cite{zeus_ffunction} and H1~\cite{h1_ffunction} collaborations;
their results were found to be 
in qualitative agreement with
those obtained in $e^+e^-$ annihilations. 
\begin{table}[ph]
\tbl{
The fractions of $c$ quarks hadronising as a particular charm hadron,
$f(c \rightarrow D, \Lambda_c)$.
The fractions are shown for the $D^+$, $D^0$, $D^+_s$
and $\Lambda_c^+$ charm ground states and for the
$D^{*+}$ state.
}
{\footnotesize
\begin{tabular}{|c|c|c|c|} \hline
& & Combined & \\
& ZEUS ($\gamma p$)~\cite{zeus_php_fb} & $e^+e^-$ data~\cite{hep-ex-9912064} & H1 (DIS)~\cite{h1_dis_fb} \\
\hline
\hline
&\phantom{~~~~~~~~~} stat.\phantom{~~} syst.\phantom{~} br.
&\phantom{~~~~}stat.$\oplus\,$syst.\phantom{~} br.
&\phantom{~~~~~~~~~~}total \\
\hline
$\fcdc$ & $0.217 \pm 0.014\phantom{~}^{+0.013\, +0.014}_{-0.005\, -0.016}$  &
$0.226\phantom{~} \pm 0.010\phantom{~~} ^{+0.016}_{-0.014}$ & $0.203 \pm 0.026$ \\
\hline
$\fcdz$ & $0.523 \pm 0.021\phantom{~}^{+0.018\, +0.022}_{-0.017\, -0.032}$  &
$0.557\phantom{~} \pm 0.023\phantom{~~} ^{+0.014}_{-0.013}$ & $0.560 \pm 0.046$ \\
\hline
$\fcdss$ & $0.095 \pm 0.008\phantom{~}^{+0.005\, +0.026}_{-0.005\, -0.017}$ &
$0.101\phantom{~} \pm 0.009\phantom{~~} ^{+0.034}_{-0.020}$ & $0.151 \pm 0.055$ \\
\hline
$\fclc$ & $0.144 \pm 0.022\phantom{~}^{+0.013\, +0.037}_{-0.022\, -0.025}$ &
$0.076\phantom{~} \pm 0.007\phantom{~~} ^{+0.027}_{-0.016}$ & \\
\hline
\hline
$\fcds$ & $0.200 \pm 0.009\phantom{~}^{+0.008\, +0.008}_{-0.006\, -0.012}$ &
$0.238\phantom{~} \pm 0.007\phantom{~~}^{+0.003}_{-0.003}$ & $0.263 \pm 0.032$ \\
\hline
\end{tabular}\label{table1} }
\vspace*{-13pt}
\end{table}

Table~1 compares the $c$-quark fragmentation fractions measured
in $ep$ interactions at HERA by
the ZEUS~\cite{zeus_php_fb} and H1~\cite{h1_dis_fb} collaborations
with those obtained in $e^+e^-$ annihilations.
The latter values were compiled previously~\cite{hep-ex-9912064}
and updated with the recent branching ratio values~\cite{pdg04}.
The measurements performed in $e^+e^-$ and $ep$ interactions are consistent.
Measurements of the $R_{u/d}$ value in charm fragmentaion confirmed
isospin invariance~\cite{zeus_php_fb,h1_dis_fb,hep-ex-9912064,zeus_dis_fb}.
Meausurements of the strangeness-suppression factor
in charm fragmentation showed
that charmed-strange meson production is suppressed
by a factor $\approx3.9$
(similar to the suppression in bottom fragmentation).
The fraction of charged $D$ mesons produced in a vector state,
$P_{\rm v}^d$,
was found to be
$\approx 0.6$\cite{zeus_php_fb,h1_dis_fb,hep-ex-9912064,zeus_dis_fb}
in both $e^+e^-$ and $ep$ interactions. The value is significantly smaller
than that obtained in bottom fragmentation and does not agree with
the naive spin counting expectation ($0.75$).

\section{Summary}

The $b$-qaurk fragmentation function and fractions were measured
in  $e^+e^-$ annihilations, while the $c$-quark fragmenation was studied
in both $e^+e^-$ and $ep$ interactions.
Comparison of the charm fragmentation characteristics, obtained
in $e^+e^-$ and $ep$ interactions, generally supports the hypothesis that
fragmentation proceeds independently of the hard sub-process.

The fraction of charged $D$ mesons produced in a vector state,
$P_{\rm v}^d$, in charm fragmentation was found to be
$\approx 0.6$
in both $e^+e^-$ and $ep$ interactions. The value is significantly smaller
than that obtained in bottom fragmentation and does not agree with
the naive spin counting expectation ($0.75$).

\end{document}